OpenStudy (anonymous):

Can anyone help me with the following problem? I don't understand the terms "root" and "constant", or their implications. One of the roots of the equation x²+kx-6=0 is 3, and k is a constant. Which is greater?: A. value of k B. -1

OpenStudy (anonymous):

root is the solution or the values of x which satisfies the given equation constant refers to any number

OpenStudy (campbell_st):

in this case the constant is the term without a pronumeral... so the constant term is -6 root is a value of x that makes the equation equal zero....

OpenStudy (campbell_st):

substitiute x = 3 into your equation (3)^2 + k(3) - 6 = 0 then 9 + 3k - 6 =0 or 3k +3 = 0 find the value of k by solving the above equation

OpenStudy (campbell_st):

the correct term for k is that is is a coefficient of x... or the number associated with the term involving x

OpenStudy (anonymous):

Ok, thank you both. So after substituting 3 for x, I've arrived at: k=-5, which means the value of answer B is greater. Did I do this right?

OpenStudy (campbell_st):

ummmm I would have said k = -1

OpenStudy (anonymous):

Ok, let me give you my process and you tell me where I went wrong: 3²+k(3)-6=0 9+3k-6=0 subtract 6 from each side 9+3k=-6 subtract 9 from each side 3k=-15 divide each side by 3 k=-5

OpenStudy (anonymous):

OH woops! I was supposed to ADD 6...lol... ok lemme try that again...

OpenStudy (anonymous):

Okay, now I get k=-1. So many details to remember!!! :) Thank you for your help!!!

OpenStudy (campbell_st):

good luck..